OpenStudy (anonymous):

prove that the length of the perpendicular from the origin to the straight line joining the two points having coordinates ( a cos alpha, a sin alpha) and (a cos beta, a sin beta) is a cos { ( alpha + beta)/2}.

OpenStudy (anonymous):

First find equation of the line passing through ( a cos alpha, a sin alpha) and (a cos beta, a sin beta)

OpenStudy (anonymous):

vat after that

OpenStudy (shamim):

may i know ur calculated eqution for those 2 given points\[(a \cos \alpha , a \sin \alpha) , (a \cos \beta ,a \sin \beta)\]

OpenStudy (anonymous):

(y - asinα)=[(asinβ-asinα)/(acosβ-acosα)]*(x-acosα)

OpenStudy (shamim):

did u solve ur problem

OpenStudy (shamim):

if u need more help then plz ask

OpenStudy (anonymous):

yes i need help